# Euclid: Early Release Observations – The star cluster systems of the Local Group dwarf galaxies IC 10 and NGC 6822††thanks: This paper is published on behalf of the Euclid Consortium

### v1

@@ Appendix A The effects of stochasticity
Stochastic sampling of the IMF can affect the integrated colours of lower-mass clusters. Fouesneau & Lançon (2010) showed that using continuous population models to analyse low-mass clusters (around and below) often leads to age underestimation. This bias also affects mass estimates, as younger models are more luminous per unit mass than older ones. While underestimation was more common, some ages were overestimated, as clusters with an excess of luminous stars above the average predicted by continuous models tend to be very red. Additionally, Popescu & Hanson (2010) demonstrated that, at these low masses, clusters may appear much redder and older than they would if their IMF was fully sampled, primarily affecting clusters with ages less than Myr but still evident at all ages in clusters below . Offsets to bluer colours exist too, but their amplitudes are less significant.
@@
We next proceeded to examine the impact of the colour variations measured above on the inferred masses and ages. We adopt fiducial colour intervals of and , where the widths reflect the maximum uncertainties typically seen in our measurements described above. We explore the properties of low-mass clusters () computed using stochastic population models with pypegase (Fouesneau & Lançon 2010), focusing on those that fall within our adopted colour intervals. By extracting the ages of all synthetic clusters within these colour ranges, we find that the age distributions remain largely independent of -band magnitude, despite the stochastic effects. Furthermore, the spread in age estimates is comparable to the typical uncertainties from SED fitting.
@@ Appendix B Tables of photometry and properties of star clusters
In these tables, we list the results of our photometry and SED-fitting analysis of the clusters (classifications 1–4). The tables include IDs, photometry, Euclid photometry, , ages, masses, metallicities (expressed in terms of , where ) and values for each cluster. Total apparent magnitudes for V band and band are provided alongside colours to convert to the remaining bands. We provide only the first six rows here; full tables will be available electronically.
@@
| ID | Age | e_Age | Mass | e_Mass | Z/Z_0 | e_Z/Z_0 | AV | e_AV |
| ESCC-NGC6822-01 | 0.1 | 0.05 | 4.44 | 0.03 | 0.03 | 0.01 | 0.73 | 0.09 |
| ESCC-NGC6822-02 | 6.21 | 0.82 | 5.81 | 0.04 | 0.04 | 0.01 | 0.78 | 0.03 |
| ESCC-NGC6822-03 | 0.7 | 0.18 | 4.46 | 0.03 | 0.24 | 0.05 | 1.02 | 0.15 |
| ESCC-NGC6822-04 | 0.55 | 0.11 | 4.15 | 0.05 | 0.17 | 0.09 | 0.96 | 0.18 |
| ESCC-NGC6822-05 | 2.49 | 0.31 | 4.52 | 0.03 | 0.19 | 0.02 | 0.67 | 0.06 |
| ESCC-NGC6822-06 | 12.0 | 1.43 | 4.36 | 0.03 | 0.24 | 0.03 | 1.07 | 0.02 |
| ID | Age | e_Age | Mass | e_Mass | Z/Z_0 | e_Z/Z_0 | AV | e_AV |
| ESCC-IC10-01 | 7.5 | 4.54 | 5.36 | 0.24 | 0.07 | 0.02 | 2.12 | 0.14 |
| ESCC-IC10-02 | 0.07 | 0.01 | 3.23 | 0.04 | 0.06 | 0.01 | 3.61 | 0.08 |
| TG3 | 0.13 | 0.03 | 3.64 | 0.04 | 0.36 | 0.12 | 2.53 | 0.06 |
| ESCC-IC10-03 | 0.63 | 0.25 | 4.41 | 0.1 | 0.1 | 0.1 | 2.65 | 0.15 |
| ESCC-IC10-04 | 0.08 | 0.02 | 3.58 | 0.05 | 0.06 | 0.01 | 2.51 | 0.05 |
| TG6 | 0.15 | 0.07 | 3.81 | 0.07 | 0.26 | 0.16 | 2.72 | 0.16 |
@@ Appendix C Cutouts of clusters
Here we present cutouts of all objects in the final catalogues (class 1–6), including newly identified clusters, previously known candidates and previously identified candidates that do not appear as clusters in the Euclid images.

### v3

@@ Appendix A The effects of stochasticity
Stochastic sampling of the IMF can affect the integrated colours of lower-mass clusters. Fouesneau and Lançon (2010) showed that using continuous population models to analyse low-mass clusters (around and below) often leads to age underestimation. This bias also affects mass estimates, as younger models are more luminous per unit mass than older ones. While underestimation was more common, some ages were overestimated, as clusters with an excess of luminous stars above the average predicted by continuous models tend to be very red. Additionally, Popescu and Hanson (2010) demonstrated that, at these low masses, clusters may appear much redder and older than they would if their IMF was fully sampled, primarily affecting clusters with ages less than Myr but still evident at all ages in clusters below . Offsets to bluer colours exist too, but their amplitudes are less significant.
@@
We next proceeded to examine the impact of the colour variations measured above on the inferred masses and ages. We adopt fiducial colour intervals of and , where the widths reflect the maximum uncertainties typically seen in our measurements described above. We explore the properties of low-mass clusters () computed using stochastic population models with pypegase (Fouesneau and Lançon 2010), focusing on those that fall within our adopted colour intervals. By extracting the ages of all synthetic clusters within these colour ranges, we find that the age distributions remain largely independent of -band magnitude, despite the stochastic effects. Furthermore, the spread in age estimates is comparable to the typical uncertainties from SED fitting.
To test the robustness of BAGPIPES in producing cluster property uncertainties commensurate with the photometric uncertainties, we drew 100 synthetic photometric realisations from the observed magnitudes and photometric uncertainties of each cluster in IC 10, and re-fit each realisation with BAGPIPES. We found agreement between the standard deviation of the resulting fits for each cluster and the single-fit posterior uncertainties. Additionally, we investigated whether the discrepancies between our age measurements and those reported in L15 (shown in Fig. 9) could arise from stochastic effects. To do this, we examined whether the age offsets between our results and those of L15 correlate with the standard deviations derived from the above 100 realisations per cluster. No correlations were apparent.
@@ Appendix B Tables of photometry and properties of star clusters
In these tables, we list the results of our photometry and SED-fitting analysis of the clusters (classifications 1–4). The tables include IDs, photometry, Euclid photometry, in pc, ages in Gyrs, for both initial mass formed and living stellar mass, metallicities (expressed in terms of , where ) and values for each cluster. Total apparent magnitudes for V band and band are provided alongside colours to convert to the remaining bands. We provide only the first six rows here; full tables are available at the CDS.
@@
| ID | Age | e_Age | Mass | e_Mass | Living Mass | e_Living Mass | Z/Z_0 | e_Z/Z_0 | AV | e_AV |
| ESCC-NGC6822-01 | 0.1 | 0.05 | 4.44 | 0.03 | 4.30 | 0.02 | 0.03 | 0.01 | 0.73 | 0.09 |
| ESCC-NGC6822-02 | 6.21 | 0.82 | 5.81 | 0.04 | 5.45 | 0.03 | 0.04 | 0.01 | 0.78 | 0.03 |
| ESCC-NGC6822-03 | 0.7 | 0.18 | 4.46 | 0.03 | 4.24 | 0.02 | 0.24 | 0.05 | 1.02 | 0.15 |
| ESCC-NGC6822-04 | 0.55 | 0.11 | 4.15 | 0.05 | 3.95 | 0.04 | 0.17 | 0.09 | 0.96 | 0.18 |
| ESCC-NGC6822-05 | 2.49 | 0.31 | 4.52 | 0.03 | 4.23 | 0.04 | 0.19 | 0.02 | 0.67 | 0.06 |
| ESCC-NGC6822-07 | 8.17 | 1.29 | 5.73 | 0.04 | 5.37 | 0.03 | 0.03 | 0.01 | 0.64 | 0.02 |
| ID | Age | e_Age | Mass | e_Mass | Living Mass | e_Living Mass | Z/Z_0 | e_Z/Z_0 | AV | e_AV |
| ESCC-IC10-01 | 7.5 | 4.54 | 5.36 | 0.24 | 4.85 | 0.22 | 0.07 | 0.02 | 2.12 | 0.14 |
| ESCC-IC10-02 | 0.07 | 0.01 | 3.23 | 0.04 | 3.05 | 0.02 | 0.06 | 0.01 | 3.61 | 0.08 |
| TG3 | 0.13 | 0.03 | 3.64 | 0.04 | 3.39 | 0.02 | 0.36 | 0.12 | 2.53 | 0.06 |
| ESCC-IC10-03 | 0.63 | 0.25 | 4.41 | 0.1 | 4.08 | 0.04 | 0.1 | 0.1 | 2.65 | 0.15 |
| ESCC-IC10-04 | 0.08 | 0.02 | 3.58 | 0.05 | 3.42 | 0.02 | 0.06 | 0.01 | 2.51 | 0.05 |
| TG6 | 0.15 | 0.07 | 3.81 | 0.07 | 3.58 | 0.04 | 0.26 | 0.16 | 2.72 | 0.16 |
@@ Appendix C Cutouts of clusters
Here we present cutouts of all objects in the final catalogues (class 1–6), including newly-identified clusters, previously-known candidates and previously-identified candidates that do not appear as clusters in the Euclid images.